OpenStudy (anonymous):

The height of a projectile shot up into the air is given by the formula h = -16t2 + 144t. Find the time t required for the projectile to return to its starting point. t= sec

hero (hero):

-16t^2+144 = 0 -t(16t -144)= 0 16t - 144 = 0 16t = 144 t = 9s

hero (hero):

You don't have to find a vertex to solve this.